# First-order and Berezinskii-Kosterlitz-Thouless phase transitions in two-dimensional generalized XY models

### v1

@@ III.2 Gibbs Measures and First Order Transitions
The Gibbs Measures method is a powerful tool for studying first-order phase transitions. For every Hamiltonian $H^{\Lambda}_{\epsilon}(\phi,\theta)$, there exists a Gibbs measure $\mu^{\Lambda}(d\phi,d\theta)$. If this measure is not unique, the system exhibits a first-order transition. A solid background on Gibbs measures and phase transitions can be found in [24, 25]. One of the early contributions using Gibbs Measures was made by Dobrushin and Shlosman [26], where they established a theorem stating that every state of Gibbs on a two-dimensional lattice with a continuous, inferiorly bounded, translationally invariant and rapidly decreasing potential is always an invariant measure under the action of a symmetry group G.
@@
While these conditions may appear restrictive, they are not impossible to satisfy. Indeed, the Heisenberg model meets these conditions, and this theorem essentially extends the Mermin-Wagner theorem, wherein invariance to the potential means the absence of continuous symmetry breaking in two-dimensional systems. The possibility of phase transitions in two-dimensional models with continuous symmetries was further explored by Shlosman in another paper[27], utilizing two powerful methods: The chessboard estimate and the reflection positivity [28, 29, 30, 31]. Shlosman demonstrated that obtaining two phases for isotropic short-range interaction models with continuous symmetry is possible, provided the temperature is low enough. However, all phases remain invariant under an action called G in the configuration space, meaning there is a likelihood of a first-order phase transition without breaking symmetry. As a practical example, Van Enter and Shlosman[10] proved this for the Domany model[6].
@@
Our focus here lies on the results obtained by Van Enter et al.[15] for the XY generalized model. They employed a square-ditch approximation, where only spins within the same ditch can interact. The present implies that a small interval $[\pi/2-\epsilon,\pi/2+\epsilon]$, is defined for the polar angle $\theta$, where the ditch parameter $\epsilon$ is a function of the $q$-parameter, i.e. $\epsilon(q)$. In addition, $\epsilon(q)$ must decay at least with $1/\sqrt{q}$. This approximation implies that two spins interact only when they are in the same ditch. The aforementioned allows them, by utilizing the reflection positivity and the chessboard estimate, to obtain an estimation of the partition function $Z^{\Lambda}$, which is given by,
$Z^{\Lambda}\geq(C_{1}\epsilon exp(2C_{2}\beta))^{|\Lambda|}\quad,$ (3)
@@
where $\beta$ is the temperature, and they assume that $C_{1}=C_{2}\approx\cos(\pi/20)\approx 1$. To demonstrate the possibility of a first-order transition, they devised another partition function. This function considers a configuration divided into two regions: an ”ordered” region where the central site and its neighbors all reside within the ditch and a ”disordered” region where the central spin and its neighbors are not in the same ditch. They integrated over all configurations compatible with this contour, resulting in a partition function with a universal contour,$Z^{\Lambda}_{uc}$ given by:
@@
In essence, if $\epsilon$ is less than 1, $Z^{\Lambda}$ dominates the fraction, resulting in a suppression of the contours between the ordered and disordered sites, leading to a completely disordered system. For $\epsilon$ that exceeds 1, the partition function $Z^{\Lambda}_{uc}$ dominates this fraction and we have a ”ordered” system. Thus, it is possible to exist a temperature region where the two partition functions are comparable, implying the non-uniqueness of the Gibbs measures.
In the work of Van Enter et al. [15], they did not estimate the critical temperature or the critical parameter q where the first-order transition appears. Here, we define a function for $\epsilon$ that preserves the system’s interactions on average as a function of the parameter $q$. This function is given by a combination of gamma functions as follows:
@@
Applying the conditions for a first-order phase transition ($\epsilon\leq 1$), we obtained the critical parameter $q_{c}\geq 2.28851$. Since $q$ is an integer, the smallest critical parameter is equal to $q_{c}=3$, which coincides with the value obtained through simulations. The previous equation also allows us to estimate the transition temperature for any $q$ value. With this equation in hand, we can again divide eq.(4) by eq.(3) and, through a simple algebraic manipulation, as follows, obtain:
@@
q FSS $\chi$ FSS $\Upsilon$ BKT $U_{L}$ First Order $U_{L}$ Gibbs Measures
@@
$\frac{Z^{\Lambda}_{uc}}{Z^{\Lambda}}\leq\frac{|(2\epsilon)^{\frac{3}{4}}exp(\beta)|^{|\Lambda|}}{|C_{1}\epsilon exp(2C_{2}\beta)|^{|\Lambda|}}\leq 1$ (7)
@@
Thus, we can estimate the critical temperature at which the first-order phase transition occurs by letting $C_{1}$ and $C_{2}$ be free parameters, expecting them to be lower than $\cos(45^{\circ})$ in accordance with the analytical results of Van Enter et al. The obtained results are shown in Table 1 and are consistent with the transition temperature obtained by histogram and Binder Fourth Cumulant. The angles obtained for both $C_{1}$ and $C_{2}$ were close to $\sim 40^{\circ}$ regardless of the parameter $q$; a value higher than that was obtained in the simulations considering the entire lattice, which was approximately $\sim 52^{\circ}$ in the transition temperature. Notably, the value of $q$ doesn’t seem to affect the azimutal angle value.
@@
In this study, we employed Monte Carlo simulations and analytical analysis to investigate the 2D generalized XY model. Our findings reveal not only characteristics of a first-order phase transition for $q\geq 3$, but also the presence of three distinct regions characterized by the existence of phase transitions. In the first region ($q\leq 3$), a BKT regime is observed where the energy and the vortex density vary continuously, consistent with the usual XY model. In the second region ($63\leq q<$ 6), a coexistence of first-order transition and BKT transition occurs, but at different temperatures, with the first-order transition temperature always higher than the BKT transition. In the third region ($q\geq 6$), the first-order transition persists, but the analyses for the critical temperature of the BKT transition do not provide conclusive results. This suggests a possible change in the universality of the system and the potential extinction of the BKT phase transition.
@@
Another important result is the impact of the vortex density on the phase transitions of the system. To illustrate, consider the histogram for $q=6$ at the transition temperature $T_{c}=0.6212$ and $L=96$. The two different phases of the system are represented by figures 4c and 4a. Figure 4a represents an ”Ordered Phase” with a lower vortex density, while figure 4c is a ”Disordered Phase” with a higher vortex density. The proliferation of vortices is identified as the cause of the first-order phase transition. To connect this result with the findings of Van Enter et al., consider the ”disordered” state, where two spins are in different ditches. The weak interaction between the spins increases the energy of the system, leading to the proliferation of vortices as shown in 4c. Conversely, when the spins belong to the same ditch, the energy decreases rapidly, resulting in a decrease in the number of vortices. Therefore, the first-order transition from an ”ordered” state to a ”disordered state” is caused by a proliferation of vortices in the system. It is essential to note that the quotation marks around ”ordered” and ”disordered” emphasize the absence of real order in the system. One phase is merely more ordered than the other, and there is no actual order in the system. Additionally, the magnetization remains null in agreement with the Theorem of Mermin-Wagner and Shlosmann, making it a first-order transition without symmetry breaking.
@@
In summary, our study offers a comprehensive understanding of the properties and phase transition behavior of the generalized XY model. The results obtained clearly demonstrate the occurrence of first-order phase transitions and elucidate their underlying mechanism. The identified classes of phase transitions align with the analytical results, and the mechanism of this new phase transition is attributed to the proliferation of vortices in the system. The next step is to explore whether other models exhibiting first-order phase transitions, such as the Domany model, can be explained in a similar manner.
@@ Acknowledgements.
The authors thank financial support from CAPES (Finance Code 001), CNPq, FAPEMIG, INCT/CNPq - Spintrônica e Nanoestruturas Magnéticas Avançadas (INCT-SpinNanoMag), and Rede Mineira de Nanomagnetismo/FAPEMIG.- Brazil. R.J.C.L. acknowledges A. Tirelli for the valuable discussions and M. Capone for all the support in developing and finalizing this work. R.J.C.L. acknowledges the financial support of the National Recovery and Resilience Plan PNRR MUR, PE0000023-NQSTI, financed by the European Union - Next Generation EU and the MUR Italian National Centre for High-Performance Computing, Big Data and Quantum Computing, grant number CN00000013, as well as the financial support from Simon’s Foundation.
Appendix A Binder’s Fourth Cumulant
@@
As a example, consider the parameter $q=3$ as in Figure 5a. We can observe the two transitions, a region where the cumulant of binder assumes negative values with a minimum at $T=0.68(2)$ for $L=128$ represents a first order transition, and a crossing point in $T=0.65(7)$ supposedly represents a BKT phase transition.
For $q=1$ to $q=5$ we can observe the BKT transition in the point where the curves intersect, however, for $q=6$ and higher the curves overlaps in more then one point, so its not possible anymore to estimate the critical temperature this way. Its a reliable technique, but different methods for determining the $T_{c}$ was considered.
Appendix B finite size scaling of magnetic susceptibility
(a) $q=3$
(b) $q=6$
(c) q=6 Binder Cumulant
Figure 6: Effects of finite-size scaling of helicity modulus for two different values of $q$ is observed in (a) and (b). With the increase of $q$, the helicity jump becomes inclined and starts at lower temperatures.(c) show FFS of the helicity modulus, for q=1 to 5, the point where the two lines cross is the critical temperature expected. 1.
@@
the exponent $\eta$ describes the long-distance behavior of in-plane correlations below $T_{c}$. For the XY model, the critical temperature is reached when $\eta=1/4$. Initially, we assume that $\eta$ is valid for any value of $q$, and the assumption is tested by the quality of the produced scaling. Thus, $\chi/L^{7/4}$ is plotted against $t$ for various system sizes. The common crossing point of the curves yields the critical temperature. For $q=1$, the expected result for the critical temperature of $T_{c}=0.69(9)$ is obtained. With an increase in the parameter $q$, the critical temperature decreases. For example, in Figure 5e, the critical temperature obtained was $t=0.65(5)$, and for $q=6$, where determining the critical temperature by this method becomes challenging, a region of crossing points is observed, as depicted in Figure 5f. This behavior persists for higher values of $q$. The interpretation drawn is that the assumption of the value of the scale parameter is no longer valid, indicating a change in the order of the phase transition.
Appendix C Helicity Modulus and Finite-Size Scalling

### v2

@@ III.2 Gibbs Measures and First-Order Transitions
The Gibbs Measures method is a powerful tool for studying first-order transitions. For every Hamiltonian $H^{\Lambda}_{\epsilon}(\phi,\theta)$, there exists a Gibbs measure $\mu^{\Lambda}(d\phi,d\theta)$. If this measure is not unique, the system exhibits a first-order transition. A solid background on Gibbs measures and phase transitions can be found in [24, 25]. One of the early contributions using Gibbs Measures was made by Dobrushin and Shlosman [26], where they established a theorem stating that every state of Gibbs on a two-dimensional lattice with a continuous, inferiorly bounded, translationally invariant and rapidly decreasing potential is always an invariant measure under the action of a symmetry group G.
@@
While these conditions may appear restrictive, they are not impossible to satisfy. Indeed, the Heisenberg model meets these conditions, and this theorem essentially extends the Mermin-Wagner theorem, wherein invariance to the potential means the absence of continuous symmetry breaking in two-dimensional systems. The possibility of phase transitions in two-dimensional models with continuous symmetries was further explored by Shlosman in another paper[27], utilizing two powerful methods: The chessboard estimate and the reflection positivity [28, 29, 30, 31]. Shlosman demonstrated that obtaining two phases for isotropic short-range interaction models with continuous symmetry is possible, provided the temperature is low enough. However, all phases remain invariant under an action called $G$ in the configuration space, meaning that there is certainty a first-order phase transition without breaking symmetry. As a another example, Van Enter and Shlosman[10] proved this for the Domany model[6].
@@
Our focus here lies on the results obtained by Van Enter et al.[15] for the generalized XY model. They employed a square-ditch approximation, where only spins within the same ditch can interact. The present implies that a small interval $[\pi/2-\epsilon,\pi/2+\epsilon]$, is defined for the polar angle $\theta$, where the ditch parameter $\epsilon$ is a function of the $q$-parameter, i.e. $\epsilon(q)$. In addition, $\epsilon(q)$ must decay at least with $1/\sqrt{q}$. This approach implies that two spins interact only when they are in the same ditch. The aforementioned allows them, by utilizing the reflection positivity and the chessboard estimate, to obtain an estimation of the partition function $Z^{\Lambda}$, which is given by,
$Z^{\Lambda}\geq(C_{1}\epsilon\exp(2C_{2}\beta))^{|\Lambda|}\quad.$ (3)
@@
where $\beta$ is the temperature, and they assume that $C_{1}=C_{2}\approx\cos(\pi/20)\approx 1$. To demonstrate the possibility of a first-order phase transition, they devised another partition function. This function considers a configuration divided into two regions: an ”ordered” region where the central site and its neighbors all reside within the ditch and a ”disordered” region where the central spin and its neighbors are not in the same ditch. They integrated over all configurations compatible with this contour, resulting in a partition function with a universal contour,$Z^{\Lambda}_{uc}$ given by:
@@
In essence, if $\epsilon$ is less than 1, $Z^{\Lambda}$ dominates the fraction, resulting in a suppression of the contours between the ordered and disordered sites, leading to a completely disordered system. If $\epsilon$ exceeds 1, the partition function $Z^{\Lambda}_{uc}$ dominates this fraction and we have a ”ordered” system. Thus, it is possible to exist a temperature region where the two partition functions are comparable, implying the non-uniqueness of the Gibbs measures.
In the work of Van Enter et al. [15], they did not estimate the critical temperature or the critical parameter q where the first-order transition appears. Here, we define a function for $\epsilon$ that preserves the system’s interactions on average as a function of the parameter $q$. This function is given by a combination of gamma functions as follows: (for a detailed derivation, see Appendix A)
@@
Applying the conditions for a first-order phase transition ($\epsilon\leq 1$), we obtained the critical parameter $q_{c}\geq 2.28851$. Note that $q$ is a measure for the narrowness of the ditch which can vary continuously. However, since we are considering only integer values of $q$, the smallest integer critical parameter would be $q_{c}=3$, which coincides with the value obtained through simulations. The previous equation also allows us to estimate the transition temperature for any $q$ value. With this equation in hand, we can again divide eq.(4) by eq.(3) and, through a simple algebraic manipulation, as follows, we get:
@@
q FSS $\chi$ FSS $\Upsilon$ BKT $U_{L}$ First-Order $U_{L}$ Histogram Gibbs Measures
@@
$\frac{Z^{\Lambda}_{uc}}{Z^{\Lambda}}\leq\frac{|(2\epsilon)^{\frac{3}{4}}\exp(\beta)|^{|\Lambda|}}{|C_{1}\epsilon\exp(2C_{2}\beta)|^{|\Lambda|}}\leq 1$ (7)
@@
Thus, we can estimate the critical temperature at which the first-order phase transition occurs by letting $C_{1}$ and $C_{2}$ be free parameters, expecting them to be lower than $\cos(45^{\circ})$ in agreement with the analytical results of Van Enter et al. The results are shown in Table 1 and are consistent with the transition temperature obtained by histogram and Binder Fourth Cumulant. The angles obtained for both $C_{1}$ and $C_{2}$ were close to $\sim 40^{\circ}$ regardless of the parameter $q$, a value lower than that was obtained in the simulations considering the entire lattice, which was approximately $\sim 52^{\circ}$ in the transition temperature. Notably, the value of $q$ doesn’t seem to affect the azimuthal angle value.
@@
In this study, we have employed Monte Carlo simulations and analytical analysis to investigate the 2D generalized XY model. Our findings reveal not only characteristics of a first-order phase transition for $q\geq 3$, but also the presence of three distinct regions characterized by the existence of phase transitions. In the first region ($q\leq 3$), a BKT regime is observed where the energy and the vortex density vary continuously, consistent with the usual XY model. In the second region ($63\leq q<$ 6), a coexistence of first-order transition and BKT transition occurs, but at different temperatures, with the first-order transition temperature always higher than the BKT transition. In the third region ($q\geq 6$), the first-order transition persists, but the analyses for the critical temperature of the BKT transition do not provide conclusive results. This suggests a possible change in the universality of the system and the potential extinction of the BKT phase transition.
@@
Another important result is the impact of the vortex density on the phase transitions of the system. To illustrate, consider the histogram for $q=6$ at the transition temperature $T_{c}=0.6212$ and $L=96$ (figure 4b). The two different phases of the system are represented by figures 4c and 4a. Figure 4a represents an ”Ordered Phase” with a lower vortex density, while figure 4c is a ”Disordered Phase” with a higher vortex density. The narrowness of the ditch causes an entropy jump [32], leading to the proliferation of vortices in the system. This is identified as the cause of the first-order phase transition. To connect this result with the findings of Van Enter et al., consider the ”disordered” state, where two spins are in different ditches. The weak interaction between the spins increases the energy of the system, leading to the proliferation of vortices as shown in 4c. Conversely, when the spins belong to the same ditch, the energy decreases rapidly, resulting in a decrease in the number of vortices. Therefore, the first-order transition from an ”ordered” state to a ”disordered state” is caused by a proliferation of vortices in the system. It is essential to note that the quotation marks around ”ordered” and ”disordered” emphasize the absence of real order in the system. One phase is merely more ordered than the other, and there is no actual order in the system. Additionally, the magnetization remains null in agreement with the Theorem of Mermin-Wagner and Van Enter-Shlosman rigorous proof, making it a disordered-disordered [33] first-order transition without symmetry breaking.
@@
In summary, our study offers a comprehensive understanding of the properties and phase transition behavior of the generalized XY model. The results obtained confirms the occurrence of first-order phase transitions and elucidate their underlying mechanism. The identified classes of phase transitions align with the analytical results, and the mechanism of this new phase transition is attributed to the proliferation of vortices in the system. The next step is to explore whether other models exhibiting first-order phase transitions, such as the Domany model, can be explained similarly.
@@ Acknowledgements.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) - Brazil - Finance Code 001. R.J.C.L. acknowledges A. Tirelli for the valuable discussions and M. Capone for all the support in developing and finalizing this work. R.J.C.L. acknowledges the financial support of the National Recovery and Resilience Plan PNRR MUR, PE0000023-NQSTI, financed by the European Union - Next Generation EU and the MUR Italian National Centre for High-Performance Computing, Big Data and Quantum Computing, grant number CN00000013, as well as the financial support from Simon’s Foundation.
Appendix A Obtaining the expression for $\epsilon(q)$
Figure 5: Graph showing the original $n(\theta)$ curve (solid lines) and the respective boxcar $nrec(\theta)$ curves (dashed lines) for different $q$ values ($q=3,5,10$).
The following steps are considered to get a q-dependent expression for $\epsilon$. First, $n(\theta)$ is replaced in the original Hamiltonian by a boxcar function, $nrec(\theta)$, as shown in the fig.(5). The boxcar function is defined in terms of two Heaviside functions, as follows:
$nrec(\theta)=A\Big{(}H\Big{[}\theta-\Big{(}\frac{\pi}{2}-\epsilon(q)\Big{)}\Big{]}-H\Big{[}\theta-\Big{(}\frac{\pi}{2}+\epsilon(q)\Big{)}\Big{]}\Big{)}$ (10)
where $A$ is the height and $H[x]$ is a Heaviside function, which is equal 00 if $x<0$ and equals to $1$ otherwise.
After the previous substitution, in order to preserve the average interactions of the system, we require that both curves, $n(\theta)$ and $nrec(\theta)$, have the same area as well as the height of $nrec(\theta)$ is equal to the average height of $n(\theta)$. Then, the equality between the areas gives:
$\mathcal{A}[nrec(\theta)]=\mathcal{A}[n(\theta)]$
$A\Big{(}2\epsilon(q)\Big{)}=\int_{0}^{\pi}\mathrm{d}\theta\sin^{q}(\theta)=\frac{\sqrt{\pi}\Gamma\left(\frac{q+1}{2}\right)}{\Gamma\left(\frac{q}{2}+1\right)}$ (11)
and, the equality of the height of $nrec(\theta)$ with the average height of $n(\theta)$ gives:
$A=\int_{0}^{\pi}\mathrm{d}\theta\left(\frac{\sqrt{\pi}\Gamma\left(\frac{q+1}{2}\right)}{\Gamma\left(\frac{q}{2}+1\right)}\right)^{-1}\sin^{2q}(\theta)$
$A=\frac{2^{-q}\sqrt{\pi}\Gamma\left(q+\frac{1}{2}\right)}{\Gamma\left(\frac{q+1}{2}\right)^{2}}$ (12)
Substituting Eq.(12) in Eq.(11), we get the equation for $\epsilon(q)$, given by:
$2\epsilon(q)\frac{2^{-q}\sqrt{\pi}\Gamma\left(q+\frac{1}{2}\right)}{\Gamma\left(\frac{q+1}{2}\right)^{2}}=\frac{\sqrt{\pi}\Gamma\left(\frac{q+1}{2}\right)}{\Gamma\left(\frac{q}{2}+1\right)}$
$\epsilon(q)=\frac{\sqrt{\pi}}{2}\left(\frac{2^{-q}\sqrt{\pi}\Gamma\left(q+\frac{1}{2}\right)}{\Gamma\left(\frac{q+1}{2}\right)^{2}}\right)^{-1}\frac{\Gamma\left(\frac{q+1}{2}\right)}{\Gamma\left(\frac{q}{2}+1\right)}$ (13)
An essential property of the ditch width, $\epsilon(q)$, as stated by [15], is that for large values of $q$, the spins only interact effectively for very narrow ditches around $\theta=\pi/2$, whose width is of the order of $\mathcal{O}(1/\sqrt{q})$. This can be easily verified from Eq.(13) by applying Stirling’s approximation to the gamma functions assuming large values of q, from which we get:
$\epsilon(q)\xrightarrow[q\to\infty]{}\frac{\sqrt{\pi}}{\sqrt{q}}$ (14)
as expected.
Appendix B Binder’s Fourth Cumulant
(a) q=3 Binder Cumulant
(b) q=3 Binder Cumulant
(c) q=6 Binder Cumulant
(d) q=3 FSS
(e) q=3 FSS
(f) q=6 FSS
Figure 6: (a),(b) is Binder Cumulant of magnetization for $q=3$ for several lattice sizes. Both crossing points for the BKT transition and the first-order phase transition region can be observed. (c) Binder Cumulant of magnetization for $q=6$ across various lattice sizes. The crossing point for the BKT transition cannot be observed anymore, but the first-order phase transition region is still visible. (c),(d) FSS of magnetic susceptibility for $q=3$ and $q=6$. For $q=3$ it’s possible to determine one crossing point, while for $q=6$ there’s a region of crossing points.
@@ Appendix B Binder’s Fourth Cumulant
As an example, consider the parameter $q=3$ as in Figure 6a. We can observe the two transitions, a region where the cumulant of Binder assumes negative values with a minimum at $T=0.68(2)$ for $L=128$ represents a first-order transition, and a crossing point in $T=0.65(7)$ supposedly represents a BKT phase transition.
From $q=1$ to $q=5$ we can observe the BKT transition at the point where the curves intersect, however, for $q=6$ and higher the curves overlap in more than one point, so it is not possible anymore to estimate the critical temperature this way. It is a reliable technique, but different methods for determining the $T_{c}$ were considered.
Appendix C finite size scaling of magnetic susceptibility
@@
the exponent $\eta$ describes the long-distance behavior of in-plane correlations below $T_{c}$. For the XY model, the critical temperature is reached when $\eta=1/4$. Initially, we assume that $\eta$ is valid for any value of $q$, and the assumption is tested by the quality of the produced scaling. Thus, $\chi/L^{7/4}$ is plotted against $t$ for various system sizes. The common crossing point of the curves yields the critical temperature. For $q=1$, the expected result for the critical temperature of $T_{c}=0.69(9)$ is obtained. With an increase in the parameter $q$, the critical temperature decreases. For example, in Figure 6f, the critical temperature obtained was $t=0.65(5)$, and for $q=6$, where determining the critical temperature by this method becomes challenging, a region of crossing points is observed, as depicted in Figure 6f. This behavior persists for higher values of $q$. The interpretation drawn is that the assumption of the value of the scale parameter is no longer valid, indicating a change in the order of the phase transition.
(a) $q=3$
(b) $q=6$
(c) q=6 Binder Cumulant
Figure 7: Effects of finite-size scaling of helicity modulus for two different values of $q$ is observed in (a) and (b). With the increase of $q$, the helicity jump becomes inclined and starts at lower temperatures. (c) show the FSS of the helicity modulus, from $q=1$ to $q=6$, the point where the two lines cross is the expected critical temperature.
Appendix D Helicity Modulus and Finite-Size Scalling